Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - Fisherman and Sophomore

Expert replies
by Resurgent » Sun Aug 23, 2009 8:44 am
In a high school debating team consisting of 2 freshmen, 2 sophomores, 2 juniors, and 2 seniors, two students are selected to represent the school at the state debating championship. The rules stipulate that the representatives must be from different grades, but otherwise the 2 representatives are to be chosen by lottery. What is the probability that the students selected will consist one freshman and one sophomore?

A. 1/16
B. 1/8
C. 1/7
D. 1/6
E. 1/4

OA: D
Join the discussion
Source: — Problem Solving |

Re: Probability - Fisherman and Sophomore

by dtweah » Sun Aug 23, 2009 1:10 pm
Resurgent wrote:In a high school debating team consisting of 2 freshmen, 2 sophomores, 2 juniors, and 2 seniors, two students are selected to represent the school at the state debating championship. The rules stipulate that the representatives must be from different grades, but otherwise the 2 representatives are to be chosen by lottery. What is the probability that the students selected will consist one freshman and one sophomore?

A. 1/16
B. 1/8
C. 1/7
D. 1/6
E. 1/4

OA: D
Let F be the event that the first student selected is Freshman. And S be the event the 2nd student is Sophomore. These two are equally likely so we will multiply our result by 2. We know that

P(S|F) = P( S and F)/P(F). We are conditioning on the fact that a freshman was chosen first. Probability of choosing a Freshman or a sophomore first is 2/8 =1/4
P( S and F)= P(F) x P(S|F)
= 1/4 x P(S|F)
This is the tricky part. If a Freshman has already been taken then we are left with 7 students over all. But the restriction says No other freshman can be chosen so the remaining freshman cannot be included in the sample space in the calculating P(S|F). So the prob that the second student is a Sophomore is 2/6 =1/3
P( S and F)= 1/4 x 1/3 =1/12

If we reverse roles we get the same value 1/12
1/12 + 1/12 = 1/6

This is an intriguing and interesting question.
Join the discussion

by ranell » Sun Aug 23, 2009 5:38 pm
(2*2)/(7*8/2) = 1/7
Join the discussion

by tom4lax » Mon Aug 24, 2009 5:15 am
F F So So J J Se Se

Choosing either Freshmen first = 2*6 (two freshmen to choose from and six remaining possibilities)

Repeat for each class gives us 48. However, need to divide by 2 so that we dont double count. (picking freshmen first and then sophmore is the same as the reverse).

So a total of 24 possibilities.

Next step is to find F and So pairings, so distinguish b/t the 4.
F1 & So1
F1 & So2
F2 & So1
F2 & So2

4/24 = 1/6
Join the discussion

by the_hustler » Mon Aug 24, 2009 9:55 am
are you sure the answer is not (1/7)?
Join the discussion

D

by enniguy » Mon Aug 24, 2009 10:49 am
Consider them 4 team with 2 people in each of them.
Now, we need to select 2 people, each from one of these 4 teams so,
4C2 = 6.

So, 6 combination exists with 2 people from different teams and only 1 combination of Fresher and Sophomore is required.

So, effectively, 1/6
Join the discussion

by rochy81 » Mon Aug 24, 2009 12:11 pm
The way I solved it was:

There's a total of 8 students to choose from. First lets choose one freshman or sophomore. There's a total of 4 freshmen and sophomores. So the probability that you can choose either one freshman or sophomore out of a total of 8 students is 4/8. Lets say we effectively get a freshman. That leaves us with 6 students left that we can choose from because we got our freshman. Now the probability that we'll choose 1 sophomore out of 6 students is 2/6. Since we're looking for one freshman AND one sophomore, we have to multiply the two probability leaving us with:

(4/8)*(2/6)=8/48 which equals 1/6.

Hope that makes sense
Join the discussion